OpenStudy (anonymous):

More cards! Consider again a standard deck of 52 cards (13 in each of 4 suits). Two cards are dealt in succession (meaning no replacement). a) What is the probability that the first card is a queen? b) Given that the first card was a queen, what is the probability that the second card is a seven? c) What is the probability of being dealt a queen first followed by a seven? d) What is the probability that both cards are greater than 7 (assuming that the ace is considered “high” or greater than 7)?

OpenStudy (anonymous):

(a) how many queens are there in the deck? well, one per suit i.e. \(4\) so the probability is \(4/52=1/13\) (as there are 52 equally-likely possibilities for the 1st card) (b) if the first card is a queen we've left only \(51\) cards however we know there are still \(4\) seven cards in there; the probability is therefore \(7/51\). note that this is higher than \(7/52\) reflecting the fact that knowing more information about the deck changes probabilities

OpenStudy (anonymous):

(c) this reflects Bayes' rule i.e. \(P(A\cap B)=P(A|B)P(B)\) so the probability is therefore \(1/13\cdot7/51=7/663\)

OpenStudy (anonymous):

I will leave (d) as an exercise